# The externally definable Ramsey property and fixed points on type spaces

### v3

@@ Abstract.
The first author is supported by Narodowe Centrum Nauki, Poland, grant 2016/22/E/ST1/00450, Israel Science Foundation grant number 555/21 and Israel Science Foundation grant number 665/20.
@@ 1. Introduction
The current paper will take a perspective more inspired by classical structural Ramsey theory: we consider Ramsey properties of structures, rather than theories, and we allow all embeddings. We are concerned with ultrahomogenoeus structures, and do not assume $\aleph_{0}$-saturation. Our central aim is to investigate the externally definable Ramsey property, where we restrict attention to colourings of a structure $M$ that are externally definable: informally, where there exists a tuple in some $N\succeq M$ such that each colour set is defined by a formula with parameters from the tuple (see Definition 3.2). (This will resemble the definitions of externally definable colouring and externally definable elementary embedding Ramsey property (EDEERP) found in [16, Definition 3.1 and Definition 3.11].) We work entirely in the context of countable relational languages.
@@
We first show (in Section 3.1) that the externally definable Ramsey property satisfies similar basic properties to those of the classical Ramsey property, with the notable exception that the equivalence between the Ramsey property for an ultrahomogeneous structure and that for its age is only shown to hold for colourings externally defined by a finite set of formulae (Proposition 3.6). (The authors believe this latter result and its use in the proofs in this paper form a new approach.) We use these properties to prove that, under the assumption of QE, if we wish to show the externally definable Ramsey property for an ultrahomogeneous structure, it suffices to check externally definable colourings of substructures of size less than the maximum arity of the language (see Proposition 3.13 - this resembles [16, Proposition 3.17], though our proof techniques differ, and as we consider structures rather than theories, we do not assume that the structure is $\aleph_{0}$-saturated). In Section 6 we then apply Proposition 3.13 to a range of examples, in an attempt to answer the first question above. A particular class of new unsaturated examples is given by lexicographic products of structures, in Section 7.
@@
    if $\operatorname{Aut}(M)$ has FPT, then $M$ has the externally definable Ramsey property;
@@
    if $\operatorname{Age}(M)$ is countable, $M$ has the externally definable Ramsey property iff $\operatorname{Aut}(M)$ has FPT.
This provides an analogue to the fundamental result of the KPT correspondence, [14, Theorem 4.5]. In the case that $M$ is $\aleph_{0}$-saturated, Theorem 5.1 follows from [16, Theorem 3.14] (personal communication, K. Krupiński). The novelty here is that Theorem 5.1 applies to all ultrahomogeneous $M$ with countable age – in particular, all Fraïssé structures, including those that are unsaturated. The dynamical characterisation of the externally definable Ramsey property as FPT is also new.
@@
The authors would also like to note that in [28], certain definable Ramsey properties for Fraïssé structures are also considered from a structural Ramsey perspective, and one particular definable Ramsey property is shown to be equivalent to having fixed points on every subflow of the Roelcke compactification of the automorphism group of the structure (see [28, Definition 1.3 and Theorem 1.8]). The relationship of these results to the results of the present paper is currently unclear (it would be interesting to know if there is any overlap in the $\aleph_{0}$-categorical case).
@@ 2.1. Ultrahomogeneity; Fraïssé theory
We assume the reader is familiar with the classical Fraïssé theory, a clear exposition of which may be found in [11, Chapter 7], and here we simply recall Fraïssé’s theorem while establishing some terminology and notation.
@@
  • if $M$ is an ultrahomogeneous structure, then $M$ has the EP;
  • if $M$ is countable and has the EP, then $M$ is ultrahomogeneous;
@@ 2.2. Quantifier elimination and $\aleph_{0}$-categoricity
We briefly review some background on quantifier elimination and $\aleph_{0}$-categoricity in the context of ultrahomogeneous structures. We assume the reader is familiar with the Ryll-Nardzewski theorem (see, for example, [19, Theorem 4.4.1]).
@@
This result follows immediately from the Ryll-Nardzewski theorem, but can also be proved directly (see [7, Proposition 1.10]).
@@
This is a corollary of the Ryll-Nardzewski theorem. A proof can be found in [7, Theorem 1.15]. By combining Lemmas 2.4 and 2.5, we obtain the following:
@@ 2.3. Ramsey theory; indivisibility
Let $\bar{a}\in\mathcal{K}$. We say that $\mathcal{K}$ has $\bar{a}$-Ramsey if for all $k\in\mathbb{N}_{+}$, for all $B\in\mathcal{K}$, there exists $C\in\mathcal{K}$ with $C\to(B)^{\bar{a}}_{k}$.
We say $\mathcal{K}$ has point-Ramsey or vertex-Ramsey if $\mathcal{K}$ has $\bar{a}$-Ramsey for all $\bar{a}\in\mathcal{K}$ with $|\bar{a}|=1$. (Usually we will only use this terminology when all structures in $\mathcal{K}$ of size $1$ are isomorphic.)
If $\mathcal{K}$ has $\bar{a}$-Ramsey for all $\bar{a}\in\mathcal{K}$, then we say that $\mathcal{K}$ has the Ramsey property. (This is also referred to as the embedding Ramsey property.)
@@
It suffices to check $\bar{a}$-Ramsey for colourings with only two colours (the proof is via a simple induction - see [2, Lemma 2.5]):
@@
If for all $B\in\mathcal{K}$ there exists $C\in\mathcal{K}$ with $C\to(B)^{\bar{a}}_{2}$, then $\mathcal{K}$ has $\bar{a}$-Ramsey.
@@
Let $M$ be a structure. Let $\bar{a}\in\operatorname{Age}(M)$. If for all $k\in\mathbb{N}_{+}$, for all $B\in\operatorname{Age}(M)$, we have that $M\to(B)^{\bar{a}}_{k}$, then we say that $M$ has $\bar{a}$-Ramsey.
The following lemma shows that to check the Ramsey property on a Fraïssé structure, it is equivalent to check the Ramsey property on its age. This is usually referred to as the “compactness principle” (see [10, Section 1.5] or [2, Proposition 2.8]).
@@
We now consider a number of indivisibility properties (see [6]), which will be useful in proving point-Ramsey for particular examples.
@@
Let $M$ be a transitive structure (i.e. $\operatorname{Aut}(M)\curvearrowright M$ is transitive). Then $M$ is age-indivisible iff $M$ has point-Ramsey.
@@
:absent:$\Rightarrow$:⇒ : Immediate. :absent:$\Leftarrow$:⇐ : Let $(S,T)$ be a partition of $M$. Let $\chi:M\to 2$ be the indicator function of $T$. Write $M=\bigcup_{i<\omega}B_{i}$ as the union of an increasing chain $B_{0}\subseteq B_{1}\subseteq\cdots$ of finite substructures of $M$. Then, as $M$ has point-Ramsey, for each $i<\omega$, $M$ contains a $\chi$-monochromatic copy of $B_{i}$. By the pigeonhole principle, there exists a colour $r\in\{0,1\}$ such that there are infinitely many $i$ with a monochromatic copy of $B_{i}$ in $M$ of colour $r$, and so $\chi^{-1}(r)$ contains a copy of each element of $\operatorname{Age}(M)$. ∎
@@ 3.1. Externally definable colourings
We say that $M$ has $\bar{a}$-Ramsey for $\Gamma$ if we have that for all $B\in\operatorname{Age}(M)$, $M\to_{\Gamma}(B)^{\bar{a}}_{k}$.
@@
again adapting the usual Erdős-Rado arrow notation. (Note that here $\Gamma$ is a set of colourings of $M$.)
We say that $\operatorname{Age}(M)$ has $\bar{a}$-Ramsey for $\Gamma$ if, for each $B\in\operatorname{Age}(M)$, there exists $C\in\operatorname{Age}(M)$ such that $C\to_{\Gamma}(B)^{\bar{a}}_{k}$.
@@
We note that the definition of an externally definable colouring resembles that of [16, Definition 3.1(b)] (see in particular [16, Remark 3.3]), except that we allow all embeddings in $\binom{M}{\bar{a}}$ rather than just elementary ones.
@@
Let $\bar{e}\in N\succeq M$ be an $|\bar{x}|$-tuple in an elementary extension $N$ of $M$ externally defining $\chi$ by the formulae $\phi_{0}(\bar{x},\bar{y}),\cdots,\phi_{k-1}(\bar{x},\bar{y})$. Let $p(\bar{x})=\operatorname{tp}_{N}(\bar{e}/M)$. Then $g\cdot p(\bar{x})$ is the type of a tuple externally defining $g\cdot\chi$ by the formulae $\phi_{0}(\bar{x},\bar{y}),\cdots,\phi_{k-1}(\bar{x},\bar{y})$. (This follows easily by compactness.) ∎
@@
We now establish a key correspondence between externally definable Ramsey properties of an ultrahomogeneous structure and externally definable Ramsey properties of its age, which we shall use frequently in the remainder of this article (this approach is new, to the best of our knowledge). Note that the below lemma is a more restricted result than the analogue in the standard Ramsey theory (Lemma 2.16), and specifically only applies to finite sets of formulae.
@@
Let $V_{\Phi}:=\operatorname{ext}(\Phi,\binom{M}{\bar{a}},k)$. Then as $\Phi$ is finite, we have that $V_{\Phi}$ is a finite union of closed sets, and hence closed. For $C\subseteq_{\text{fin.\!}}M$, let $W_{C}$ denote the set of colourings $\chi:\binom{M}{\bar{a}}\to k$ such that $C$ contains no $\chi$-monochromatic copy of $B$. It is easy to check that $W_{C}$ is closed. For $C\subseteq M$, let $V_{C}=V_{\Phi}\cap W_{C}$. By assumption (using Lemma 3.5, which does not require countability of $M$), each $V_{C}$ is non-empty. Therefore the collection of closed sets $\{V_{C}\mid C\subseteq_{\text{fin.\!}}M\}$ has the finite intersection property, and as $k^{\binom{M}{\bar{a}}}$ is compact, we have that $\bigcap_{C\subseteq_{\text{fin.\!}}M}V_{C}$ is non-empty. Thus we have $M\not\to_{\operatorname{ext}(\Phi)}(B)^{\bar{a}}_{k}$. ∎
@@
Let $M$ be an ultrahomogeneous structure. Let $\bar{a}\in\operatorname{Age}(M)$. Let $\Phi$ be a set of formulae (where free variables in $\bar{y}$ are interpreted as $\bar{a}$).
Suppose that $M$ has $\bar{a}$-Ramsey for all $2$-colourings externally defined by a formula from $\Phi$. Then $M$ has $\bar{a}$-Ramsey for all $2$-colourings externally defined by a formula consisting of a Boolean combination of formulae from $\Phi$.
@@
For $s\in\mathbb{N}$, let $\Psi_{s}=\{\psi\in\Psi:l(\psi)\leq s\}$. Using induction on $s\in\mathbb{N}$, we will show that $M$ has $\bar{a}$-Ramsey for each $2$-colouring externally defined by a formula in $\Psi_{s}$.
The base case is given by the assumption that $M$ has $\bar{a}$-Ramsey for all $2$-colourings externally defined by a formula from $\Phi$, as $\Psi_{0}=\Phi$.
Assume that $M$ has $\bar{a}$-Ramsey for $2$-colourings externally defined by a formula from $\Psi_{s}$, and let $\chi:\binom{M}{\bar{a}}\to 2$ be a $2$-colouring externally defined by $\bar{e}\in N\succeq M$ and a formula $\psi(\bar{x},\bar{y})\in\Psi_{s+1}$ of complexity $s+1$. We must have one of the following:
@@
In the first case, by the induction assumption the colouring $\chi_{1}:\binom{M}{\bar{a}}\to 2$ externally defined by $\bar{e}\in N\succeq M$ and $\psi_{1}(\bar{x},\bar{y})$ has $\bar{a}$-Ramsey, and as $\chi=1-\chi_{1}$, thus $\chi$ has $\bar{a}$-Ramsey.
In the second case, as $\psi_{1}(\bar{x},\bar{y})$, $\psi_{2}(\bar{x},\bar{y})$ are of complexity $\leq s$, by the induction assumption $M$ has $\bar{a}$-Ramsey for colourings externally defined by $\psi_{1}(\bar{x},\bar{y})$ and for colourings externally defined by $\psi_{2}(\bar{x},\bar{y})$. Let $\chi_{1}:\binom{M}{\bar{a}}\to 2$ be the colouring externally defined by $\bar{e}\in N\succeq M$ and $\psi_{1}(\bar{x},\bar{y})$, and let $\chi_{2}:\binom{M}{\bar{a}}\to 2$ be the colouring externally defined by $\bar{e}\in N\succeq M$ and $\psi_{2}(\bar{x},\bar{y})$. Take $B\in\operatorname{Age}(M)$. Applying Proposition 3.6 to the set $\{\psi_{2}(\bar{x},\bar{y})\}$, there is $C\in\operatorname{Age}(M)$ with $C\to_{\chi_{2}}(B)^{\bar{a}}_{2}$, i.e. for any copy $C^{\prime}$ of $C$ in $M$, $C^{\prime}$ contains a $\chi_{2}$-monochromatic copy of $B$. As $M$ has $\bar{a}$-Ramsey for $\chi_{1}$, there is a $\chi_{1}$-monochromatic copy $C^{\prime}\subseteq M$ of $C$. $C^{\prime}$ contains a $\chi_{2}$-monochromatic copy $B^{\prime}$ of $B$, and so $B^{\prime}$ is monochromatic in $\chi_{1}$ and $\chi_{2}$. Therefore $B^{\prime}$ is monochromatic in $\chi$. ∎

### v4

@@ Abstract.
The first author is supported by Israel Science Foundation grant number 555/21, Israel Science Foundation grant number 665/20, and Narodowe Centrum Nauki, Poland, grant 2016/22/E/ST1/00450.
@@ 1. Introduction
The current paper will take a perspective more inspired by classical structural Ramsey theory: we consider Ramsey properties of structures, rather than theories, and we allow all embeddings. We are concerned with ultrahomogeneous structures, and do not assume $\aleph_{0}$-saturation. Our central aim is to investigate the externally definable Ramsey property, where we restrict attention to colourings of a structure $M$ that are externally definable: informally, where there exists a tuple in some $N\succeq M$ such that each colour set is defined by a formula with parameters from the tuple (see Definition 3.2). (This will resemble the definitions of externally definable colouring and externally definable elementary embedding Ramsey property (EDEERP) found in [15, Definition 3.1 and Definition 3.11].) We work entirely in the context of countable relational languages.
@@
We first show (in Section 3.1) that the externally definable Ramsey property satisfies similar basic properties to those of the classical Ramsey property, with the notable exception that the equivalence between the Ramsey property for an ultrahomogeneous structure and that for its age is only shown to hold for colourings externally defined by a finite set of formulae (Proposition 3.6). (The authors believe this latter result and its use in the proofs in this paper form a new approach.) We use these properties to prove that, under the assumption of QE, if we wish to show the externally definable Ramsey property for an ultrahomogeneous structure, it suffices to check externally definable colourings of substructures of size less than the maximum arity of the language (see Proposition 3.13 - this resembles [15, Proposition 3.17], though our proof techniques differ, and as we consider structures rather than theories, we do not assume that the structure is $\aleph_{0}$-saturated). In Section 6 we then apply Proposition 3.13 to a range of examples, in an attempt to answer the first question above. A particular class of new unsaturated examples is given by lexicographic products of structures, in Section 7.
@@
    if $M$ has FPT, then $M$ has the externally definable Ramsey property;
@@
    if $\operatorname{Age}(M)$ is countable, then $M$ has the externally definable Ramsey property iff $M$ has FPT.
This provides an analogue to the fundamental result of the KPT correspondence, [13, Theorem 4.5]. In the case that $M$ is $\aleph_{0}$-saturated, Theorem 5.1 follows from [15, Theorem 3.14] (personal communication, K. Krupiński). The novelty here is that Theorem 5.1 applies to all ultrahomogeneous $M$ with countable age – in particular, all Fraïssé structures, including those that are unsaturated. The dynamical characterisation of the externally definable Ramsey property as FPT is also new.
@@
The authors would also like to note that in [27], certain definable Ramsey properties for Fraïssé structures are also considered from a structural Ramsey perspective, and one particular definable Ramsey property is shown to be equivalent to having fixed points on every subflow of the Roelcke compactification of the automorphism group of the structure (see [27, Definition 1.3 and Theorem 1.8]). The relationship of these results to the results of the present paper is currently unclear (it would be interesting to know if there is any overlap in the $\aleph_{0}$-categorical case).
@@ 2.1. Ultrahomogeneity; Fraïssé theory
We assume the reader is familiar with the classical Fraïssé theory, a clear exposition of which may be found in [10, Chapter 7], and here we simply recall Fraïssé’s theorem while establishing some terminology and notation.
@@
  • if $M$ is an ultrahomogeneous structure, then $M$ has the extension property (EP);
  • if $M$ is countable and has the extension property, then $M$ is ultrahomogeneous;
@@ 2.2. Quantifier elimination and $\aleph_{0}$-categoricity
We briefly review some background on quantifier elimination and $\aleph_{0}$-categoricity in the context of ultrahomogeneous structures. We assume the reader is familiar with the Ryll-Nardzewski theorem (see, for example, [18, Theorem 4.4.1]).
@@
This result follows immediately from the Ryll-Nardzewski theorem, but can also be proved directly (see [6, Proposition 1.10]).
@@
This is a corollary of the Ryll-Nardzewski theorem. A proof can be found in [6, Theorem 1.15]. By combining Lemmas 2.4 and 2.5, we obtain the following:
@@ 2.3. Ramsey theory; indivisibility
Let $\bar{a}\in\mathcal{K}$. We say that $\mathcal{K}$ has the $\bar{a}$-Ramsey property ($\bar{a}$-RP) if for all $k\in\mathbb{N}_{+}$, for all $B\in\mathcal{K}$, there exists $C\in\mathcal{K}$ with $C\to(B)^{\bar{a}}_{k}$.
We say $\mathcal{K}$ has the point-Ramsey property or vertex-Ramsey property if $\mathcal{K}$ has the $\bar{a}$-RP for all $\bar{a}\in\mathcal{K}$ with $|\bar{a}|=1$. (Usually we will only use this terminology when all structures in $\mathcal{K}$ of size $1$ are isomorphic.)
If $\mathcal{K}$ has the $\bar{a}$-Ramsey property for all $\bar{a}\in\mathcal{K}$, then we say that $\mathcal{K}$ has the Ramsey property. (This is also referred to as the embedding Ramsey property.)
@@
It suffices to check the $\bar{a}$-Ramsey property for colourings with only two colours (the proof is via a simple induction - see [2, Lemma 2.5]):
@@
If for all $B\in\mathcal{K}$ there exists $C\in\mathcal{K}$ with $C\to(B)^{\bar{a}}_{2}$, then $\mathcal{K}$ has the $\bar{a}$-RP.
@@
Let $M$ be a structure. Let $\bar{a}\in\operatorname{Age}(M)$. If for all $k\in\mathbb{N}_{+}$, for all $B\in\operatorname{Age}(M)$, we have that $M\to(B)^{\bar{a}}_{k}$, then we say that $M$ has the $\bar{a}$-Ramsey property.
As in Definition 2.12, we say that $M$ has the point-Ramsey property if $M$ has the $\bar{a}$-Ramsey property for all $\bar{a}\in\operatorname{Age}(M)$ with $|\bar{a}|=1$, and we say that $M$ has the Ramsey property if $M$ has the $\bar{a}$-Ramsey property for all $\bar{a}\in\operatorname{Age}(M)$.
The following lemma shows that to check the Ramsey property on a Fraïssé structure, it is equivalent to check the Ramsey property on its age. This is usually referred to as the “compactness principle” (see [9, Section 1.5] or [2, Proposition 2.8]).
@@
We now consider a number of indivisibility properties (see [5]), which will be useful in proving the point-Ramsey property for particular examples.
@@
Let $M$ be a countable transitive structure (i.e. $\operatorname{Aut}(M)\curvearrowright M$ is transitive). Then $M$ is age-indivisible iff $M$ has the point-Ramsey property.
@@
:absent:$\Rightarrow$:⇒ : Immediate. :absent:$\Leftarrow$:⇐ : Let $(S,T)$ be a partition of $M$. Let $\chi:M\to 2$ be the indicator function of $T$. Write $M=\bigcup_{i<\omega}B_{i}$ as the union of an increasing chain $B_{0}\subseteq B_{1}\subseteq\cdots$ of finite substructures of $M$. Then, as $M$ has the point-Ramsey property, for each $i<\omega$, $M$ contains a $\chi$-monochromatic copy of $B_{i}$. By the pigeonhole principle, there exists a colour $r\in\{0,1\}$ such that there are infinitely many $i$ with a monochromatic copy of $B_{i}$ in $M$ of colour $r$, and so $\chi^{-1}(r)$ contains a copy of each element of $\operatorname{Age}(M)$. ∎
@@ 3.1. Externally definable colourings
We say that $M$ has the $\bar{a}$-Ramsey property for $\Gamma$ ($\bar{a}$-RP for $\Gamma$) if we have that for all $B\in\operatorname{Age}(M)$, $M\to_{\Gamma}(B)^{\bar{a}}_{k}$.
@@
again adapting the usual Erdős-Rado arrow notation. (Note that here $\Gamma$ is a set of colourings of $M$, and we consider the colourings induced by $\Gamma$ on each copy of $C$ in $M$. See Lemma 3.5.)
We say that $\operatorname{Age}(M)$ has the $\bar{a}$-Ramsey property for $\Gamma$ ($\bar{a}$-RP for $\Gamma$) if, for each $B\in\operatorname{Age}(M)$, there exists $C\in\operatorname{Age}(M)$ such that $C\to_{\Gamma}(B)^{\bar{a}}_{k}$.
@@
We note that the definition of an externally definable colouring resembles that of [15, Definition 3.1(b)] (see in particular [15, Remark 3.3]), except that we allow all embeddings in $\binom{M}{\bar{a}}$ rather than just elementary ones.
@@
If $M\to_{\operatorname{ext}}(B)^{\bar{a}}_{k}$ for all $B\in\operatorname{Age}(M)$, we say that $M$ has the externally definable $\bar{a}$-Ramsey property (externally definable $\bar{a}$-RP). If $M$ has the externally definable $\bar{a}$-Ramsey property for all $\bar{a}\in\operatorname{Age}(M)$, we say that $M$ has the externally definable Ramsey property.
@@
Let $\bar{e}\in N\succeq M$ be an $|\bar{x}|$-tuple in an elementary extension $N$ of $M$ externally defining $\chi$ by the formulae $\phi_{0}(\bar{x},\bar{y}),\cdots,\phi_{k-1}(\bar{x},\bar{y})$. Let $p(\bar{x})=\operatorname{tp}_{N}(\bar{e}/M)$. Then $g\cdot p(\bar{x})$ is the type of a tuple externally defining $g\cdot\chi$ by the formulae $\phi_{0}(\bar{x},\bar{y}),\cdots,\phi_{k-1}(\bar{x},\bar{y})$. ∎
@@
We now establish a key correspondence between externally definable Ramsey properties of an ultrahomogeneous structure and externally definable Ramsey properties of its age, which we shall use frequently in the remainder of this article (this approach is new, to the best of our knowledge). Note that the below proposition is a more restricted result than the analogue in the standard Ramsey theory (Lemma 2.16), and specifically only applies to finite sets of formulae.
@@
Let $V_{\Phi}:=\operatorname{ext}(\Phi,\binom{M}{\bar{a}},k)$. Then as $\Phi$ is finite, we have that $V_{\Phi}$ is a finite union of closed sets, and hence closed. For $C\subseteq_{\text{fin.\!}}M$, let $W_{C}$ denote the set of colourings $\chi:\binom{M}{\bar{a}}\to k$ such that $C$ contains no $\chi$-monochromatic copy of $B$. It is easy to check that each $W_{C}$ is closed, and that for $C_{0},C_{1}\subseteq_{\text{fin.\!}}M$, we have $W_{C_{0}\cup C_{1}}\subseteq W_{C_{0}}\cap W_{C_{1}}$. For $C\subseteq_{\text{fin.\!}}M$, let $V_{C}=V_{\Phi}\cap W_{C}$. By assumption (using Lemma 3.5), each $V_{C}$ is non-empty. Therefore the collection of closed sets $\{V_{C}\mid C\subseteq_{\text{fin.\!}}M\}$ has the finite intersection property, and as $k^{\binom{M}{\bar{a}}}$ is compact, we have that $\bigcap_{C\subseteq_{\text{fin.\!}}M}V_{C}$ is non-empty. Thus we have $M\not\to_{\operatorname{ext}(\Phi)}(B)^{\bar{a}}_{k}$. ∎
@@
Let $M$ be an ultrahomogeneous structure. Let $\bar{a}\in\operatorname{Age}(M)$. Let $\Phi$ be a set of formulae (where free variables in $\bar{y}$ correspond to $\bar{a}$).
Suppose that $M$ has the $\bar{a}$-RP for all $2$-colourings externally defined by a formula from $\Phi$. Then $M$ has the $\bar{a}$-RP for all $2$-colourings externally defined by a formula consisting of a Boolean combination of formulae from $\Phi$.
@@
For $s\in\mathbb{N}$, let $\Psi_{s}=\{\psi\in\Psi:l(\psi)\leq s\}$. Using induction on $s\in\mathbb{N}$, we will show that $M$ has the $\bar{a}$-RP for each $2$-colouring externally defined by a formula in $\Psi_{s}$.
The base case is given by the assumption that $M$ has the $\bar{a}$-RP for all $2$-colourings externally defined by a formula from $\Phi$, as $\Psi_{0}=\Phi$.
Assume that $M$ has the $\bar{a}$-RP for $2$-colourings externally defined by a formula from $\Psi_{s}$, and let $\chi:\binom{M}{\bar{a}}\to 2$ be a $2$-colouring externally defined by $\bar{e}\in N\succeq M$ and a formula $\psi(\bar{x},\bar{y})\in\Psi_{s+1}$ of complexity $s+1$. We must have one of the following:
@@
For the first case, take the colouring $\chi_{1}:\binom{M}{\bar{a}}\to 2$ externally defined by $\bar{e}\in N\succeq M$ and $\psi_{1}(\bar{x},\bar{y})$. Then $\chi=1-\chi_{1}$, and so by the induction assumption $M$ has the $\bar{a}$-RP for $\chi$.
In the second case, as $\psi_{1}(\bar{x},\bar{y})$, $\psi_{2}(\bar{x},\bar{y})$ are of complexity $\leq s$, by the induction assumption $M$ has the $\bar{a}$-RP for colourings externally defined by $\psi_{1}(\bar{x},\bar{y})$ and for colourings externally defined by $\psi_{2}(\bar{x},\bar{y})$. Let $\chi_{1}:\binom{M}{\bar{a}}\to 2$ be the colouring externally defined by $\bar{e}\in N\succeq M$ and $\psi_{1}(\bar{x},\bar{y})$, and let $\chi_{2}:\binom{M}{\bar{a}}\to 2$ be the colouring externally defined by $\bar{e}\in N\succeq M$ and $\psi_{2}(\bar{x},\bar{y})$. Take $B\in\operatorname{Age}(M)$. Applying Proposition 3.6 to the set $\{\psi_{2}(\bar{x},\bar{y})\}$, there is $C\in\operatorname{Age}(M)$ with $C\to_{\chi_{2}}(B)^{\bar{a}}_{2}$, i.e. for any copy $C^{\prime}$ of $C$ in $M$, $C^{\prime}$ contains a $\chi_{2}$-monochromatic copy of $B$. As $M$ has the $\bar{a}$-RP for $\chi_{1}$, there is a $\chi_{1}$-monochromatic copy $C^{\prime}\subseteq M$ of $C$. $C^{\prime}$ contains a $\chi_{2}$-monochromatic copy $B^{\prime}$ of $B$, and so $B^{\prime}$ is monochromatic in $\chi_{1}$ and $\chi_{2}$. Therefore $B^{\prime}$ is monochromatic in $\chi$. ∎